# Soft Theorems in Chern-Simons Matter Theories

### v3

@@ Abstract
We describe tree level soft theorems for Chern-Simons QED and QCD in 4+1 dimensions. Soft theorems have been studied in various context in the past. The universal behavior of the leading and subleading soft theorems is understood as a consequence of the local on-shell invariance of the scattering amplitudes. Chern-Simons terms are gauge invariant only up to a boundary term. Hence, they make for an interesting case to study the relationship between gauge invariance and soft theorems. In this paper, we show that the subleading soft theorems in QED and QCD get corrected when Chern-Simons terms are introduced and also derive those corrections. We start with a discussion on soft theorems as a consequence of gauge invariance. Then we move on to derive the Chern-Simons vertices. Finally, we calculate the correction to soft theorems due to the addition of Chern-Simons terms for both QED and QCD.
@@
Throughout this paper, I have worked in natural units with $\hbar=c=1$. I have also used the “mostly minus” signature $\left(+---\right)$ for the space-time metric. $\epsilon^{\mu\nu\rho\sigma}$ refers to the Levi-Civita symbol normalized as $\epsilon^{0123}=1$. $\mu$, $\nu$ denote the Lorentz indices while $\alpha$, $\beta$ denote the indices of the representation of the gauge group. The latter will often be omitted to tidy up the equations, in which case the order of multiplication is important. Lorentz indices are raised and lowered with the Minkowski metric $\eta_{\mu\nu}$. Whenever mentioned, the $\gamma^{\mu}$ denote the Dirac matrices satisfying the Clifford algebra $\left\{\gamma^{\mu},\gamma^{\nu}\right\}=2\eta^{\mu\nu}$. We use the Feynman slash notation where $\cancel{b}=b^{\mu}\gamma_{\mu}$ for any four vector $b^{\mu}$. $A_{\mu}$, $a_{\mu}$ are spin-1 gauge-fields in Lorentz gauge with the latter (with small $a$) denoting a soft gauge mode. $F_{\mu\nu}$ is the field strength tensor. We will be working with SU($N$) gauge theories for which $a_{\mu}$ etc. are matrices in the representation. They can be expanded in terms of the generators of the gauge group $a_{\mu}\left(x\right)=a_{a\mu}\left(x\right)T_{a}$. The following types of shorthand for index contractions will be present,
@@
We will use the superscript $T$ to denote the matrix transpose in the representation indices $\alpha$, $\beta$. Einstein summation convention is used for all index contractions. Therefore,
@@ 1 Introduction
Soft theorems are statements about scattering amplitudes of processes involving external states with low energy (soft) gauge bosons. They relate the amplitudes with external soft gauge bosons to those without them via a soft operator. This soft operator is given as an expansion in the soft momenta of the external gauge bosons. The leading term in the expansion has a pole divergence in the soft momenta Gell-Mann:1954wra; Low:1954kd; Low:1958sn; Burnett:1967km; Weinberg:1964ew; Weinberg:1965nx. The leading and subleading terms in the soft operator have a universal form which is understood to be a consequence of local on-shell gauge invariance of the scattering amplitudes Low:1954kd; Low:1958sn; Weinberg:1964ew; Burnett:1967km; Bern:2014vva. This naturally leads to the following question: does this statement apply to gauge theories with Chern-Simons terms as well? Unlike the usual gauge invariant terms in the QED or QCD action, the Chern-Simons terms are gauge invariant only up to a boundary term Chern:1974ft; Dunne:1998qy. Therefore, they do not modify the Ward identity corresponding to small gauge transformations. Naturally then it is interesting to ask whether they modify the universal form of the leading or subleading soft theorems. In this paper, we answer this question. We calculate how the soft theorems get corrected at the tree level when we introduce a Chern-Simons term in QED and QCD.
@@
Chern-Simons terms are only defined in odd space-time dimensions. They are usually studied in 2+1 space-time dimensions. We are instead working in 4+1 dimensions because we want to treat the Chern-Simons term perturbatively about the usual gauge theory of QED or QCD. [^1] In D=2+1, the Chern-Simons term goes like $\sim\mathbb{A}d\mathbb{A}+\mathbb{A}^{3}$ which dominates over the Yang Mills action in the infrared. This not the case in higher dimensions where the Chern-Simons term becomes less relevant in the infrared. Hence, we consider a gauge field coupled to matter in $D=4+1$,
@@
$S_{gauge}=\int d^{4+1}x\left\{-\frac{1}{4g^{2}}F^{2}-\frac{1}{2g^{2}}\left(\partial.A\right)^{2}+\mathcal{L}_{matter}+\mathcal{L}_{int}\right\},$ (1)
where $A_{\mu}$ is the gauge field and $F_{\mu\nu}$ is the appropriate field strength tensor. In the case of QED, the gauge field is abelian and hence we add the abelian Chern-Simons term given by,
$\mathcal{S}_{CS}$ $=\kappa\int\mathbb{A}\left(d\mathbb{A}\right)^{2}=\kappa\int d^{4+1}x\,\epsilon^{\mu\nu\rho\sigma\delta}A_{\mu}\partial_{\nu}A_{\rho}\partial_{\sigma}A_{\delta}.$ (2)
where we have introduced the gauge one form $\mathbb{A}=A_{\mu}dx^{\mu}$ and the product is to be understood as the wedge product $\wedge$. For QCD, we need to add the non-abelian Chern-Simons form in 4+1 dimensions,
$S_{CS}=\kappa\int\mbox{Tr}\left[\mathbb{A}\left(d\mathbb{A}\right)^{2}+c_{4}\mathbb{A}^{3}d\mathbb{A}+c_{5}\mathbb{A}^{5}\right],$ (3)
@@
When $\left\{c_{4},c_{5}\right\}=\left\{\frac{3}{2},\frac{3}{5}\right\}$, $\mathcal{S}_{CS}$ becomes a Chern-Simons term Miskovic:2005di and is gauge invariant up to a boundary term [^2] For non-abelian Chern-Simons term, under large gauge transformations the action is also shifted by a constant which is referred to as the winding number Dunne:1998qy. This does not affect the path integral since for appropriate discrete values of $\kappa$, the winding number is $2\pi$ times an integer and hence leaves $e^{i\mathcal{S}}$ invariant. . The total action is then given by the sum of the two,
$S=S_{gauge}+S_{CS},$
@@
In the absence of the Chern-Simons term $S_{CS}$, the subleading soft theorem (for instance, for an external soft photon with momentum $k_{\mu}$ and polarization $e_{\mu}$) takes the form,
$G_{n+1}\left(e,k;\varepsilon_{i},p_{i}\right)=\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e.p_{i}}{p_{i}.k}G_{n}^{\left(i\right)}\left(p_{i}\right)-i\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e_{\mu}k_{\nu}}{p_{i}.k}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right),$ (4)
@@
where $\varepsilon_{i},p_{i}$ refers to the polarization and energy of the external finite energy (“hard”) particles and $q_{i}$ refers to the charge of the matter coupling to the photon field. Similar expressions are there for gluons and gravitons as well. The leading term has a $\nicefrac{{1}}{{k}}$ pole divergence while the subleading term is $\mathcal{O}\left(k^{0}\right)$. The soft photon theorem at leading order in $k$ does not get infrared corrections at loop level Bern:2014vva; Bern:2014oka; He:2014bga; He:2014cra. In $D=2+1$, the subleading term gets corrected due to infrared effects at loop level, but in $D=4+1$ the subleading soft photon theorem is also valid to all loop orders. The soft gluon theorem in $D=2+1$ admits infrared corrections at loop level even at the leading order but in $D=4+1$, the theory becomes IR finite and corrections occur only at subleading order. In the case of multiple soft gauge bosons, each gauge boson brings a soft factor of divergence $\nicefrac{{1}}{{k}}$ at leading order.
@@
As mentioned before, the leading and subleading behavior in (4) is universal and can be understood as a consequence of local on-shell gauge invariance of the scattering amplitudes. This was shown for photons in Low:1958sn and for gluons in Bern:2014vva; Broedel:2014fsa. For gravitons, even the second subleading behaviour (not stated here) is universal and can be obtained from gauge invariance of the amplitudes Weinberg:1964ew; Weinberg:1965nx; Bern:2014vva; Broedel:2014fsa. Sen’s covaraintisation procedure Sen:2017nim; Sen:2017xjn; AtulBhatkar:2018kfi; Laddha:2017ygw; Chakrabarti:2017ltl utilizes gauge invariance to derive this universal behavior by giving a covariantisation prescription to calculate the 3-point vertex (see $\Gamma^{\left(3\right)}\left(e,k;p,-p-k\right)$ in figure 1) for the leading order diagram. Here, one writes an action involving only hard particle fields and covariantises it by minimally coupling the gauge field mode for the soft gauge boson. This is described in detail in section 2. Such a covariantisation prescription does not apply to Chern-Simons terms because even though the small gauge Ward identity is still valid, the Chern-Simons terms are not minimally coupled to the matter. Hence, their contribution needs to be accounted for separately in the 3-point vertex, $\Gamma^{\left(3\right)}\left(e,k;p,-p-k\right)$.
@@
Furthermore, we show that gauge invariance is a sufficient but not necessary condition for having the universal form of the leading soft theorems. For this, we will take $c_{4}$ and $c_{5}$ in (3) to be arbitrary (and hence make $\mathcal{S}_{CS}$ non gauge-invariant) and show that the universal form of the leading soft theorem still holds and only the subleading term gets corrected.
@@
Hence, ultimately the aim of this paper is to calculate how (4) gets corrected at tree level when a Chern-Simons term $\mathcal{S}_{CS}$ is added to the gauge theory.
@@
The paper is organized as follows - in section §2, we review the soft photon and gluon theorems and derive them by covariantising the hard particle action by minimally coupling the soft gauge field modes. AtulBhatkar:2018kfi; Sen:2017nim; Chakrabarti:2017ltl; Sen:2017xjn; Laddha:2017ygw demonstrate it for soft photons and gravitons. We generalize it to a generic non-abelian Yang Mills theory with SU($N$) gauge invariance and derive the soft gluon theorem. Soft gluon theorems have been previously studied in various ways in Bern:2014vva; Mao:2017tey; Broedel:2014fsa; Broedel:2014bza; Klose:2015xoa. Then in section §3, we calculate the Feynman rules for the Chern-Simons vertices in $D=4+1$. In section §4, §5 and §6 we incorporate them to obtain the Chern-Simons correction to the soft theorems. We see that the leading soft theorems remain unchanged and the Chern-Simons terms only correct the subleading soft theorems.
@@ 2 Subleading Soft Theorem from Sen’s Covariantisation Procedure
In this section we will summarize the subleading single soft theorem for photons, as given in AtulBhatkar:2018kfi using Sen’s covariantisation procedure Sen:2017nim; AtulBhatkar:2018kfi; Chakrabarti:2017ltl; Sen:2017xjn; Laddha:2017ygw and generalize it for gluons. For gluons, the amplitudes and the corresponding diagrams will be implicitly assumed to be color ordered. At loop level, the familiar soft theorems (4) get subleading infrared corrections in case of photons and leading infrared corrections in case of gluons. Therefore, we will restrict ourselves to only tree level amplitudes even in this general covariantisation procedure. We start by considering a scattering process for hard particles, with amplitude $G_{n}\left(p_{i}\right)$. We assume that the hard particles in the process are described by a single set of fields $\phi_{\alpha}$, belonging to some reducible representation of the Lorentz group. Let $\mathcal{S}_{hard}\left[\phi\right]$ denote the corresponding action for the hard particle fields. We assume that this action is invariant under the global transformation,
$\phi_{\alpha}\rightarrow\left[e^{i\theta_{a}T_{a}}\right]_{\alpha}^{\,\,\,\,\beta}\phi_{\beta}.$
@@
Here $\Gamma^{\left(3\right)}$ is the tree level 3 point vertex involving a soft boson as well as two hard particles in the external state. $G_{n}^{\left(i\right),\alpha_{i}}$ is the full $n$-point amplitude involving $n$ external hard particles with $i$th particle polarization amputated. To evaluate these diagrams we need to find the vertices. These can be obtained by taking $S_{hard}$ and covariantising it with respect to the single mode soft gauge field $\left[a_{\mu}\left(x\right)\right]_{\alpha}^{\,\,\,\,\beta}=\left[e_{\mu}\left(k\right)\right]_{\alpha}^{\,\,\,\,\,\beta}\exp\left(ik.x\right)$ to get $S_{soft}$. This involves replacing all ordinary derivatives $\partial_{\mu}$ in the action with covariant derivatives $D_{\mu}$ as follows-
$D_{\mu}\phi_{\alpha}=\partial_{\mu}\phi_{\alpha}-i\left[a_{\mu}\right]_{\alpha}^{\,\,\,\,\beta}\phi_{\beta},$ (5)
@@
To be more explicit, suppose we want to evaluate $\Gamma^{\left(3\right)}$. For this, consider the quadratic part of $S_{hard}$ which can be written as [^3] Note that we use $\phi$ to denote the field in both position space and momentum space.,
$S_{hard}^{\left(2\right)}$ $=\frac{1}{2}\int d^{D}xd^{D}y\phi_{\alpha}\left(x\right)\hat{K}^{\alpha\beta}\left(x,y\right)\phi_{\beta}\left(y\right)$ (6)
@@
$\Delta_{\alpha\beta}\left(p\right)=iK_{\alpha\beta}^{-1}\left(p\right)=\frac{\Xi_{\alpha\beta}\left(p\right)}{p^{2}-m^{2}},$ (10)
here $\Xi_{\alpha\beta}$ is the polarization sum (times $i$) for the hard particle fields. Invariance of (6) under global transformation (5) implies,
@@
$\varepsilon_{\alpha}\left(p\right)K^{\alpha\beta}\left(-p\right)=0,$ (13)
@@
$\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\Xi\left(-p\right)=2ip^{\mu}-K\left(-p\right)\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}},$ (15)
$\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}}K\left(-p\right)=2ip^{\mu}-\Xi\left(-p\right)\frac{\partial K\left(-p\right)}{\partial p_{\mu}},$ (16)
@@
Covariantising of (6) with respect to a single soft mode gauge field $a_{\mu}\left(x\right)=e_{\mu}\left(k\right)\exp\left(ik.x\right)$ using the procedure above gives the following terms in the action $S_{soft}$ contributing to the vertex $\Gamma^{\left(3\right)}$ ,
$S_{soft}^{\left(3\right)}=$ $\frac{1}{2}\int\frac{d^{D}p_{1}}{\left(2\pi\right)^{D}}\frac{d^{D}p_{2}}{\left(2\pi\right)^{D}}\left(2\pi\right)^{D}\delta^{\left(D\right)}\left(p_{1}+p_{2}+k\right)$
@@
$S_{soft}^{\left(3\right)}=$ $\frac{1}{2}\int\frac{d^{D}p_{1}}{\left(2\pi\right)^{D}}\frac{d^{D}p_{2}}{\left(2\pi\right)^{D}}\left(2\pi\right)^{D}\delta^{\left(D\right)}\left(p_{1}+p_{2}+k\right)$
@@
$\frac{\partial K}{\partial p_{\mu}}\left(p\right)=iC_{2}^{\mu}-2C_{3}^{\mu\nu}p_{\nu},$ $\hskip 10.00002pt\frac{\partial^{2}K}{\partial p_{\mu}\partial p_{\nu}}\left(p\right)=-2C_{3}^{\mu\nu}.$
@@
$D_{\mu}D_{\nu}\phi_{\alpha}=$ $\left(\left.\delta_{\alpha}\right.^{\gamma}\partial_{\mu}-ia_{\mu}{}_{\alpha}^{\,\,\,\,\gamma}\right)\left(\left.\delta_{\gamma}\right.^{\beta}\partial_{\nu}-ia_{\nu}{}_{\gamma}^{\,\,\,\,\beta}\right)\phi_{\beta}$
$=$ $\partial_{\mu}\partial_{\nu}\phi_{\alpha}-i\left(\partial_{\mu}a_{\nu}\right)_{\alpha}^{\,\,\,\,\beta}\phi_{\beta}-i\left(a_{\mu}\partial_{\nu}+a_{\nu}\partial_{\mu}\right)_{\alpha}^{\,\,\,\,\beta}\phi_{\beta}+\mathcal{O}\left(a^{2}\right),$
@@
$\left[C_{1}+C_{2}^{\mu}D_{\mu}+C_{3}^{\mu\nu}D_{\mu}D_{\nu}\right]\phi$ $\xRightarrow{\left(\text{momentum space}\right)}\left[K\left(p\right)-iC_{2}^{\mu}e_{\mu}+2C_{3}^{\mu\nu}p_{\mu}e_{\nu}+C_{3}^{\mu\nu}k_{\mu}e_{\nu}\right]\phi$

### v4

@@ Abstract
We describe tree-level soft theorems for Chern-Simons QED and QCD in 4+1 dimensions. Soft theorems have been studied in various context in the past. The universal behavior of the leading and subleading soft theorems is understood as a consequence of the local on-shell gauge invariance of the scattering amplitudes. Chern-Simons terms are gauge-invariant only up to a boundary term. Hence, they make for an interesting case to study the relationship between gauge invariance and soft theorems. In this paper, we show that the subleading soft theorems in QED and QCD get corrected when Chern-Simons terms are introduced and also derive those corrections. We start with a discussion on soft theorems as a consequence of gauge invariance. Then we move on to derive the Chern-Simons vertices. Finally, we calculate the correction to soft theorems due to the addition of Chern-Simons terms for both QED and QCD.
@@
Throughout this paper, we work in natural units with $\hbar=c=1$. We also use the “mostly minus” signature $\left(+---\right)$ for the space-time metric. $\epsilon^{\mu\nu\rho\sigma}$ refers to the Levi-Civita symbol normalized as $\epsilon^{0123}=1$. $\mu$, $\nu$ denote the Lorentz indices while $\alpha$, $\beta$ denote the indices of the representation of the gauge group. The latter is often omitted to tidy up the equations, in which case the order of multiplication is important. Lorentz indices are raised and lowered with the Minkowski metric $\eta_{\mu\nu}$. Whenever mentioned, the $\gamma^{\mu}$ denote the Dirac matrices satisfying the Clifford algebra $\left\{\gamma^{\mu},\gamma^{\nu}\right\}=2\eta^{\mu\nu}$. We use the Feynman slash notation where $\cancel{b}=b^{\mu}\gamma_{\mu}$ for any four vector $b^{\mu}$. $A_{\mu}$, $a_{\mu}$ are spin-1 gauge-fields in Lorentz gauge with the latter (with small $a$) denoting a soft gauge mode. $F_{\mu\nu}$ is the field strength tensor. We work with SU($N$) gauge theories for which $a_{\mu}$ etc. are matrices in the representation. They can be expanded in terms of the generators of the gauge group $a_{\mu}\left(x\right)=a_{a\mu}\left(x\right)T_{a}$. The following types of shorthand for index contractions will be present,
@@
We also use the superscript $T$ to denote the matrix transpose in the representation indices $\alpha$, $\beta$. Einstein summation convention is used for all index contractions. Therefore,
@@ 1 Introduction
Soft theorems are statements about scattering amplitudes of processes involving external states with low energy (soft) gauge bosons. They relate the amplitudes with external soft gauge bosons to those without them via a soft operator. This soft operator is given as an expansion in the soft momenta of the external gauge bosons. The leading term in the expansion has a pole divergence in the soft momenta [10, 15, 16, 6, 20, 21]. The leading and subleading terms in the soft operator have a universal form which is completely insensitive to the details of the process and only depends on the charge, momenta and angular momentum of the hard particles. This universal behavior is understood to be a consequence of the local on-shell gauge invariance of the scattering amplitudes [15, 16, 20, 6, 2]. This naturally leads to the following question: does this statement apply to gauge theories with Chern-Simons terms as well? Unlike the usual gauge-invariant terms in the QED or QCD action, the Chern-Simons terms are gauge-invariant only up to a boundary term [8, 9]. Therefore, they do not modify the Ward identity corresponding to small gauge transformations. Therefore, the gauge invariance of these terms is more subtle. Naturally then it is interesting to ask whether they modify the universal form of the leading or subleading soft theorems. In this paper, we answer this question. We calculate how the soft theorems get corrected at the tree-level when we introduce a Chern-Simons term in QED and QCD.
@@
Chern-Simons terms are only defined in odd space-time dimensions. They are usually studied in 2+1 space-time dimensions. We are instead working in 4+1 dimensions because we want to treat the Chern-Simons term perturbatively about the usual gauge theory of QED or QCD. [^1] In D=2+1, the Chern-Simons term goes like $\sim\mathbb{A}d\mathbb{A}+\mathbb{A}^{3}$ which dominates over the Yang-Mills action in the infrared. This not the case in higher dimensions where the Chern-Simons term becomes less relevant in the infrared. Hence, we consider a gauge field coupled to matter in $D=4+1$,
@@
$S_{\text{gauge}}=\int d^{4+1}x\left\{-\frac{1}{4g^{2}}F^{2}-\frac{1}{2g^{2}}\left(\partial.A\right)^{2}+\mathcal{L}_{\text{matter}}+\mathcal{L}_{\text{int}}\right\},$ (1)
where $A_{\mu}$ is the gauge field and $F_{\mu\nu}$ is the appropriate field strength tensor.
To this we can then add a deformation of the form,
$S_{\text{CS}}=\kappa\int\mbox{Tr}\left[\mathbb{A}\left(d\mathbb{A}\right)^{2}+c_{4}\mathbb{A}^{3}d\mathbb{A}+c_{5}\mathbb{A}^{5}\right],$ (2)
@@
In the case of QED, the gauge field is Abelian and hence the last two terms vanish. The first term by itself is then the Abelian Chern-Simons term. For QCD, the deformation becomes gauge-invariant (up to a boundary term) when $\left\{c_{4},c_{5}\right\}=\left\{\frac{3}{2},\frac{3}{5}\right\}$. Then $S_{\text{CS}}$ becomes the non-Abelian Chern-Simons term. [^2] For non-Abelian Chern-Simons term, under large gauge transformations the action is also shifted by a constant which is referred to as the winding number [9]. This does not affect the path integral since for appropriate discrete values of $\kappa$, the winding number is $2\pi$ times an integer and hence leaves $e^{\mathrm{i}S}$ invariant.
The total action is then given by the sum of the two,
$S=S_{\text{gauge}}+S_{\text{CS}},$
@@
In the absence of the term $S_{\text{CS}}$, the subleading soft theorem (for instance, for an external soft photon with momentum $k_{\mu}$ and polarization $e_{\mu}$) takes the form,
$G_{n+1}\left(e,k;\varepsilon_{i},p_{i}\right)=\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e.p_{i}}{p_{i}.k}G_{n}^{\left(i\right)}\left(p_{i}\right)-\mathrm{i}\sum_{i}\varepsilon_{i}^{T}q_{i}\frac{e_{\mu}k_{\nu}}{p_{i}.k}\left(\hat{J}_{\left(i\right)}^{\mu\nu}G_{n}^{\left(i\right)}\right)\left(p_{i}\right),$ (3)
@@
where $\varepsilon_{i},p_{i}$ refers to the polarization and energy of the external finite energy (“hard”) particles and $q_{i}$ refers to the charge of the matter coupling to the photon field. Similar expressions are there for gluons and gravitons as well. The leading term has a $\nicefrac{{1}}{{k}}$ pole divergence while the subleading term is $\mathcal{O}\left(k^{0}\right)$. The soft photon theorem at leading order in $k$ does not get infrared corrections at loop level [2, 3, 11, 12]. In $D=2+1$, the subleading term gets corrected due to infrared effects at loop level, but in $D=4+1$ the subleading soft photon theorem is also valid to all loop orders. The soft gluon theorem in $D=2+1$ admits infrared corrections at loop level even at the leading order but in $D=4+1$, the theory becomes IR finite and corrections occur only at subleading order. In the case of multiple soft gauge bosons, each gauge boson brings a soft factor of divergence $\nicefrac{{1}}{{k}}$ at leading order.
@@
As mentioned before, the leading and subleading behavior in (3) is universal and can be understood as a consequence of local on-shell gauge invariance of the scattering amplitudes. This was shown for photons in [16] and for gluons in [2, 4]. For gravitons, even the second subleading behaviour (not stated here) is universal and can be obtained from gauge invariance of the amplitudes [20, 21, 2, 4]. Sen’s covaraintisation procedure [19, 18, 1, 14, 7] utilizes gauge invariance to derive this universal behavior by giving a covariantisation prescription to calculate the 3-point vertex (see $\Gamma^{\left(3\right)}\left(e,k;p,-p-k\right)$ in figure 1) for the leading order diagram. Here, one writes an action involving only hard particle fields and covariantises it by minimally coupling the gauge field mode for the soft gauge boson. This is described in detail in section 2. Such a covariantisation prescription does not apply to Chern-Simons terms because even though the small gauge Ward identity is still valid, the Chern-Simons terms are not minimally coupled to the matter. Hence, their contribution needs to be accounted for separately in the 3-point vertex, $\Gamma^{\left(3\right)}\left(e,k;p,-p-k\right)$.
@@
Furthermore, we show that gauge invariance is a sufficient but not necessary condition for having the universal form of the leading soft theorems. For this, we will take $c_{4}$ and $c_{5}$ in (2) to be arbitrary (and hence make $S_{\text{CS}}$ non-gauge-invariant) and show that the universal form of the leading soft theorem still holds and only the subleading term gets corrected.
@@
Hence, ultimately the aim of this paper is to calculate how (3) gets corrected at tree-level when a Chern-Simons term $S_{\text{CS}}$ is added to the gauge theory.
@@
The paper is organized as follows - in section §2, we review the soft photon and gluon theorems and derive them by covariantising the hard particle action by minimally coupling the soft gauge field modes. [1, 19, 7, 18, 14] demonstrate it for soft photons and gravitons. We generalize it to a generic non-Abelian Yang-Mills theory with SU($N$) gauge invariance and derive the soft gluon theorem. Soft gluon theorems have been previously studied in various ways in [2, 17, 4, 5, 13]. Then in section §3, we calculate the Feynman rules for the Chern-Simons vertices in $D=4+1$. In section §4, §5 and §6 we incorporate them to obtain the Chern-Simons correction to the soft theorems. We see that the leading soft theorems remain unchanged and the Chern-Simons terms only correct the subleading soft theorems.
@@ 2 Subleading Soft Theorem from Sen’s Covariantisation Procedure
In this section we will summarize the subleading single soft theorem for photons, as given in [1] using Sen’s covariantisation procedure [19, 1, 7, 18, 14] and generalize it for gluons. For gluons, the amplitudes and the corresponding diagrams will be implicitly assumed to be color ordered. At loop level, the familiar soft theorems (3) get subleading infrared corrections in case of photons and leading infrared corrections in case of gluons. Therefore, we will restrict ourselves to only tree-level amplitudes even in this general covariantisation procedure. We start by considering a scattering process for hard particles, with amplitude $G_{n}\left(p_{i}\right)$. We assume that the hard particles in the process are described by a single set of fields $\phi_{\alpha}$, belonging to some reducible representation of the Lorentz group. Let $S_{\text{hard}}\left[\phi\right]$ denote the corresponding action for the hard particle fields. We assume that this action is invariant under the global transformation,
$\phi_{\alpha}\rightarrow\left[e^{\mathrm{i}\theta_{a}T_{a}}\right]_{\alpha}^{\,\,\,\,\beta}\phi_{\beta}.$
@@
Here $\Gamma^{\left(3\right)}$ is the tree-level 3-point vertex involving a soft boson as well as two hard particles in the external state. $G_{n}^{\left(i\right),\alpha_{i}}$ is the full $n$-point amplitude involving $n$ external hard particles with $i$th particle polarization amputated. To evaluate these diagrams we need to find the vertices. These can be obtained by taking $S_{\text{hard}}$ and covariantising it with respect to the single mode soft gauge field $\left[a_{\mu}\left(x\right)\right]_{\alpha}^{\,\,\,\,\beta}=\left[e_{\mu}\left(k\right)\right]_{\alpha}^{\,\,\,\,\,\beta}\exp\left(\mathrm{i}k.x\right)$ to get $S_{\text{soft}}$. This involves replacing all ordinary derivatives $\partial_{\mu}$ in the action with covariant derivatives $D_{\mu}$ as follows-
$D_{\mu}\phi_{\alpha}=\partial_{\mu}\phi_{\alpha}-\mathrm{i}\left[a_{\mu}\right]_{\alpha}^{\,\,\,\,\beta}\phi_{\beta},$ (4)
@@
To be more explicit, suppose we want to evaluate $\Gamma^{\left(3\right)}$. For this, consider the quadratic part of $S_{\text{hard}}$ which can be written as [^3] Note that we use $\phi$ to denote the field in both position space and momentum space.,
$S_{\text{hard}}^{\left(2\right)}$ $=\frac{1}{2}\int d^{D}xd^{D}y\phi_{\alpha}\left(x\right)\hat{K}^{\alpha\beta}\left(x,y\right)\phi_{\beta}\left(y\right)$ (5)
@@
$\Delta_{\alpha\beta}\left(p\right)=\mathrm{i}K_{\alpha\beta}^{-1}\left(p\right)=\frac{\Xi_{\alpha\beta}\left(p\right)}{p^{2}-m^{2}},$ (9)
here $\Xi_{\alpha\beta}$ is the polarization sum (times $\mathrm{i}$) for the hard particle fields. Invariance of (5) under global transformation (4) implies,
@@
$\varepsilon_{\alpha}\left(p\right)K^{\alpha\beta}\left(-p\right)=0.$ (12)
@@
$\frac{\partial K\left(-p\right)}{\partial p_{\mu}}\Xi\left(-p\right)=2\mathrm{i}p^{\mu}-K\left(-p\right)\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}},$ (14)
$\frac{\partial\Xi\left(-p\right)}{\partial p_{\mu}}K\left(-p\right)=2\mathrm{i}p^{\mu}-\Xi\left(-p\right)\frac{\partial K\left(-p\right)}{\partial p_{\mu}},$ (15)
@@
Covariantising of (5) with respect to a single soft mode gauge field $a_{\mu}\left(x\right)=e_{\mu}\left(k\right)\exp\left(\mathrm{i}k.x\right)$ using the procedure above gives the following terms in the action $S_{\text{soft}}$ contributing to the vertex $\Gamma^{\left(3\right)}$ ,
$S_{\text{soft}}^{\left(3\right)}=$ $\frac{1}{2}\int\frac{d^{D}p_{1}}{\left(2\pi\right)^{D}}\frac{d^{D}p_{2}}{\left(2\pi\right)^{D}}\left(2\pi\right)^{D}\delta^{\left(D\right)}\left(p_{1}+p_{2}+k\right)$
@@
$S_{\text{soft}}^{\left(3\right)}=$ $\frac{1}{2}\int\frac{d^{D}p_{1}}{\left(2\pi\right)^{D}}\frac{d^{D}p_{2}}{\left(2\pi\right)^{D}}\left(2\pi\right)^{D}\delta^{\left(D\right)}\left(p_{1}+p_{2}+k\right)$
@@
$\frac{\partial K}{\partial p_{\mu}}\left(p\right)=\mathrm{i}C_{2}^{\mu}-2C_{3}^{\mu\nu}p_{\nu},$ $\hskip 10.00002pt\frac{\partial^{2}K}{\partial p_{\mu}\partial p_{\nu}}\left(p\right)=-2C_{3}^{\mu\nu}.$
@@
$D_{\mu}D_{\nu}\phi_{\alpha}=$ $\left(\left.\delta_{\alpha}\right.^{\gamma}\partial_{\mu}-\mathrm{i}\,a_{\mu}{}_{\alpha}^{\,\,\,\,\gamma}\right)\left(\left.\delta_{\gamma}\right.^{\beta}\partial_{\nu}-\mathrm{i}\,a_{\nu}{}_{\gamma}^{\,\,\,\,\beta}\right)\phi_{\beta}$
$=$ $\partial_{\mu}\partial_{\nu}\phi_{\alpha}-\mathrm{i}\,\left(\partial_{\mu}a_{\nu}\right)_{\alpha}^{\,\,\,\,\beta}\phi_{\beta}-\mathrm{i}\,\left(a_{\mu}\partial_{\nu}+a_{\nu}\partial_{\mu}\right)_{\alpha}^{\,\,\,\,\beta}\phi_{\beta}+\mathcal{O}\left(a^{2}\right),$
@@
$\left[C_{1}+C_{2}^{\mu}D_{\mu}+C_{3}^{\mu\nu}D_{\mu}D_{\nu}\right]\phi$ $\xRightarrow{\left(\text{momentum space}\right)}\left[K\left(p\right)-\mathrm{i}\,C_{2}^{\mu}e_{\mu}+2C_{3}^{\mu\nu}p_{\mu}e_{\nu}+C_{3}^{\mu\nu}k_{\mu}e_{\nu}\right]\phi$